OpenStudy (anonymous):

If a loaded truck that can accelerate at 2 m/s2 loses its load and has one-third of its original mass, what acceleration can it attain if the same driving force acts on it?

OpenStudy (anonymous):

F=ma (original mass) F=ma (1/3 original mass) Because you want the same driving force (same F) for both situations, you can set these two equations = to each other. ma (original mass) = ma (1/3 original mass) You are given the acceleration of the orginal mass, 2m/s^2. You know the second situation has 1/3 of the orginal mass. So (m)*(2m/s^2) = 1/3(m)*(a); because m is on each side they cancel so: 2m/s^2=1/3a Solve for a by multiply the other side by 3. (2m/s^2)*(3) = a = 6m/s^2 This makes sense because in order to maintain the same magnitude of F, the acceleration would need to make up for the loss of mass.

OpenStudy (anonymous):

here, 1st case, acceleration(a)= 2 m/s^2 force=f (say) mass=m (say) then, f=ma=mx2=2m......(1) 2nd case acceleration(a)=a' (say) force=f mass=m/3 then, f=ma=(m/3)xa'.......(2) from question force "f" is equal in both the cases.then, equation 1=equation 2 i.e. 2m=(m/3)xa' or,a'=(2x3)m/m=6 m/s^2 so,to act same driving force on truck acceleration must be 6 m/s^2 when mass becomes one-third of its original mass.